# Secure Digital Semantic Communications: Fundamentals, Challenges, and Opportunities

### v4

@@ II Background Knowledge of Semantic Communication
Figure 3: An overview of the security and privacy threats against digital SemCom. The red dashed lines link threats to the corresponding stages where they arise. For probabilistic modulation, demodulation is often integrated with the JSC decoder, whereas for deterministic modulation, demodulation and the JSC decoder are typically implemented as separate modules. For clarity and a unified presentation, we depict demodulation and the JSC decoder as two distinct blocks to highlight the threats targeting each stage. This overview includes threats shared by both analog and digital SemCom as well as those unique to digital SemCom. In this paper, we primarily focus on threats specific to digital SemCom and their corresponding defenses.
@@ III Security and Privacy Threats Against Digital Semantic Communication
This section characterizes the security and privacy threat landscape against digital SemCom. We first summarize threats common to analog and digital SemCom, and then examine vulnerabilities specific to digital implementations, with emphasis on finite-alphabet semantic transmission and packet-based delivery. An overview of the security and privacy threats against digital SemCom is illustrated in Fig. 3.
@@ III-B1 Attacks on Probabilistic Modulation
In the branch that targets discrete channel models such as the binary erasure channel (BEC) and binary symmetric channel (BSC), the encoder acts as a neural-network-parameterized Bernoulli generator whose outputs determine per-bit sampling probabilities. A primary threat is probability biasing, where an adversary induces small but structured perturbations to the source or intermediate semantic information, and thereby causes shifts in a subset of Bernoulli parameters. As a result, the sampled bit patterns become more likely to trigger semantic decoding errors at the receiver.
@@ III-B2 Attacks on Deterministic Modulation
In scalar quantization, an adversary can perturb selected feature elements to push them across quantization thresholds. Depending on the integer-to-bit mapping, such threshold crossings can induce multiple bit-level changes and thereby increase the likelihood of semantic decoding errors. In symbol quantization, remapping learned symbol-like outputs onto a predefined constellation introduces decision regions that can be exploited by carefully designed perturbations. Such perturbations can shift received points across region boundaries and cause systematic symbol decision errors.. In vector quantization with a shared codebook, semantic vectors are represented by codebook indices before bit mapping and conventional digital modulation. Perturbations near codeword boundaries can induce nearest-codeword confusion, and corruption of index bits can redirect decoding toward unintended codewords. Because vector quantization relies on a shared codebook, codebook integrity and consistency between endpoints are critical. Compromising the codebook through poisoning, unauthorized updates, or desynchronization can lead to systematic semantic misinterpretation.
@@ IV-B Defenses Against Attacks on Probabilistic Modulation
Attacks on probabilistic modulation mainly aim to reshape the learned conditional distribution or to disrupt sampling behavior. A practical defense objective is therefore to make the transmitter distribution less sensitive to small perturbations in the source or intermediate semantic information. It should also prevent attack-induced inflation of predictive uncertainty or overconfident concentration that can amplify semantic errors. This can be supported in two ways. First, by regularizing the probabilistic encoder so that its output distribution remains reasonably well calibrated and exhibits controlled uncertainty under channel variations. Second, by using perturbation-aware training so that the learned distribution changes smoothly when the source or the semantic information is slightly perturbed. As a result, the distribution avoids large probability-mass shifts and abrupt concentration changes. In addition, sampling-induced inconsistency in the semantic decoding outcome can be reduced by adopting controlled sampling at deployment. For example, the receiver can draw a limited number of samples and aggregate them, enforce confidence- or entropy-based constraints, or switch to more conservative selection rules when predicted uncertainty is high. These measures reduce the adversary’s ability to induce unstable task outputs through distribution manipulation while introducing extra latency only when predicted risk is high.
@@ IV-C Defenses Against Attacks on Deterministic Modulation
Attacks on deterministic modulation are enabled by the explicit quantizer-dequantizer interface as well as the resulting hard decision boundaries and fixed mapping rules. Defenses should therefore strengthen the discretization interface to reduce the likelihood that small feature perturbations change discrete outcomes, while improving the robustness of discrete representations as they traverse practical digital physical-layer procedures and packet-based delivery. Practical measures include quantizer and mapping designs with explicit safety margins around decision thresholds, as well as boundary-aware objectives that discourage operating near decision boundaries. They also include representation formats that limit the impact of localized errors, such as mappings with limited error propagation and lightweight error detection for discrete intermediates before semantic decoding.
For scalar quantization, defenses can emphasize robust bit mapping and bitstream protection. Gray-style mappings reduce multi-bit changes caused by a single threshold crossing. Interleaving or scrambling, together with error detection or correction, can mitigate localized corruption before it causes large semantic distortions in downstream semantic decoding. For symbol quantization, robustness can be improved by conservative constellation remapping with sufficient separation between decision regions, and boundary-aware training to reduce sensitivity near decision boundaries. For vector quantization, protection should secure codebook distribution and updates, validate codebook versions, and protect transmitted indices with encryption and integrity checks, since small index changes can redirect decoding to an unintended codeword and cause significant semantic deviation. Codebook integrity monitoring and authenticated updates further reduce the risk of poisoning or unauthorized modifications that manipulate codeword boundaries or endpoint consistency.
@@ IV-D Defenses Against Packet- and Protocol-Level Threats
To mitigate active attacks, digital SemCom can rely on standard packet security and protocol hardening, while accounting for the semantic setting where a small subset of packets may dominate task performance. Authenticated integrity protection defends against injection and manipulation, and anti-replay mechanisms with sequence validation mitigate replay and reordering for both payload packets and semantic-related control messages. Robustness to selective dropping can be improved through cross-packet redundancy and unequal error protection for task-critical packets. In addition, receiver-side plausibility checks can detect inconsistencies in recovered semantic information under missing or altered packets, particularly when high-impact packets are selectively removed. Fragmentation and reassembly should be validated and hardened. This prevents malformed fragments and ambiguous reassembly outcomes from corrupting semantic information or bypassing sequence validation and integrity checks. Protocol operations such as acknowledgments, retransmissions, and rate adaptation should be protected against forged feedback. This can be achieved by authenticating feedback messages and applying conservative fallback rules when the reported state is inconsistent or statistically implausible. Configuration and synchronization exchanges should be authenticated, including coding and modulation settings, semantic model versions, codebooks, and knowledge base updates. Strict version validation and rollback prevention are also needed. These measures help avoid persistent transmitter-receiver mismatch that can degrade semantic decoding quality over multiple packets.

### v6

@@ II Background Knowledge of Semantic Communication
Figure 3: An overview of the security and privacy threats against digital SemCom. The red dashed lines link threats to the corresponding stages where they arise. For probabilistic modulation, demodulation is often integrated with the JSC decoder, whereas for deterministic modulation, demodulation and the JSC decoder are typically implemented as separate modules. For clarity and a unified presentation, we depict demodulation and the JSC decoder as two distinct blocks to highlight the threats targeting each stage. This overview includes threats shared by both analog and digital SemCom as well as those unique to digital SemCom. In particular, modulation-related threats and packet-channel threats are distinctive features of digital SemCom and are generally absent in analog implementations. In this paper, we primarily focus on threats specific to digital SemCom and their corresponding defenses. Blocks with a star symbol in the upper-left corner denote threats that are also present in conventional digital communication systems but may be amplified in SemCom.
@@
While conventional digital communication also involves modulation, digital SemCom changes the modulation interface because the transmitted bits or symbols are produced from learned semantic features and optimized for a downstream task. In probabilistic modulation, the transmitter outputs a conditional distribution over a discrete alphabet and generates the transmitted sequence by sampling, so an adversary can bias the recovered semantic information by manipulating the learned distribution or the sampling process. In deterministic modulation, discretization is realized via quantization, and in vector quantization variants the shared codebook and its indices become part of the semantic representation, so boundary crossing, index corruption, or codebook manipulation can cause systematic semantic remapping and task-level degradation.
@@ III Security and Privacy Threats Against Digital Semantic Communication
This section characterizes the security and privacy threat landscape against digital SemCom. We first summarize threats common to analog and digital SemCom, and then examine vulnerabilities specific to digital implementations, with emphasis on finite-alphabet semantic transmission and packet-based delivery. An overview of the security and privacy threats against digital SemCom is illustrated in Fig. 3. To clarify the boundary with conventional digital communication systems, in Fig. 3, blocks with a star symbol in the upper-left corner denote threats that are also present in conventional digital communication systems but may be amplified in SemCom. Blocks without the star symbol denote threats that are specific to SemCom systems.
@@ III-B1 Attacks on Probabilistic Modulation
In the branch that targets discrete channel models such as the binary erasure channel (BEC) and binary symmetric channel (BSC), the encoder acts as a neural-network-parameterized Bernoulli generator whose outputs determine per-bit sampling probabilities. A primary threat is probability biasing, where an adversary induces structured perturbations to the source or intermediate semantic information, and thereby causes shifts in a subset of Bernoulli parameters. As a result, the sampled bit patterns become more likely to trigger semantic decoding errors at the receiver.
@@ III-B2 Attacks on Deterministic Modulation
In scalar quantization, an adversary can perturb selected feature elements to push them across quantization thresholds. Depending on the integer-to-bit mapping, such threshold crossings can induce multiple bit-level changes and thereby increase the likelihood of semantic decoding errors. In symbol quantization, remapping learned symbol-like outputs onto a predefined constellation introduces decision regions that can be exploited by carefully designed perturbations. Such perturbations can shift received points across region boundaries and cause symbol decision errors. In vector quantization with a shared codebook, semantic vectors are represented by codebook indices before bit mapping and conventional digital modulation. Perturbations near codeword boundaries can induce nearest-codeword confusion, and corruption of index bits can redirect decoding toward unintended codewords. Because vector quantization relies on a shared codebook, codebook integrity and consistency between endpoints are critical. Compromising the codebook through poisoning, unauthorized updates, or desynchronization can lead to systematic semantic misinterpretation.
@@ IV-B Defenses Against Attacks on Probabilistic Modulation
Attacks on probabilistic modulation mainly aim to reshape the learned conditional distribution or to disrupt sampling behavior. A practical defense objective is therefore to make the transmitter distribution less sensitive to small perturbations in the source or intermediate semantic information. It should also prevent attack-induced inflation of predictive uncertainty or overconfident concentration that can amplify semantic errors. This can be supported by regularizing the probabilistic encoder to maintain a reasonably well-calibrated output distribution and controlled uncertainty under channel variations. In particular, we can adopt perturbation-aware training to enforce smooth changes in the learned distribution under small perturbations of the source or semantic information. As a result, the distribution avoids large probability-mass shifts and abrupt concentration changes. In addition, sampling-induced inconsistency in the semantic decoding outcome can be reduced by adopting controlled sampling at deployment. For example, the receiver can draw a limited number of samples and aggregate them, enforce confidence- or entropy-based constraints, or switch to more conservative selection rules when predicted uncertainty is high. These measures reduce the adversary’s ability to induce unstable task outputs through distribution manipulation while introducing extra latency only when predicted risk is high.
@@ IV-C Defenses Against Attacks on Deterministic Modulation
Attacks on deterministic modulation are enabled by the explicit quantizer-dequantizer interface as well as the resulting hard decision boundaries and fixed mapping rules. Defenses should therefore strengthen the discretization interface to reduce the likelihood that small feature perturbations change discrete outcomes, while improving the robustness of discrete representations as they traverse practical digital physical-layer procedures and packet-based delivery. Practical measures include quantizer and mapping designs with explicit safety margins around decision thresholds, as well as boundary-aware objectives that discourage operating near decision thresholds. They also include representation formats that limit the impact of localized errors, such as mappings with limited error propagation and lightweight error detection for discrete intermediates before semantic decoding.
For scalar quantization, defenses can emphasize robust bit mapping and bitstream protection. Gray-style mappings reduce multi-bit changes caused by a single threshold crossing. Interleaving or scrambling, together with error detection or correction, can mitigate localized corruption before it causes large semantic distortions in downstream semantic decoding. For symbol quantization, robustness can be improved by conservative constellation remapping with sufficient separation between decision regions, and boundary-aware training to reduce sensitivity near decision thresholds. For vector quantization, protection should secure codebook distribution and updates, validate codebook versions, and protect transmitted indices with encryption and integrity checks, since small index changes can redirect decoding to an unintended codeword and cause significant semantic deviation. Codebook integrity monitoring and authenticated updates further reduce the risk of poisoning or unauthorized modifications that manipulate codeword boundaries or endpoint consistency. These defenses may constrain representation flexibility due to more conservative boundary designs. In addition, maintaining consistent codebook versions and authenticated updates across transmitter and receiver increases system-level coordination requirements.
@@ IV-D Defenses Against Packet- and Protocol-Level Threats
To mitigate active attacks, digital SemCom can rely on standard packet security and protocol hardening, while accounting for the semantic setting where a small subset of packets may dominate task performance. Authenticated integrity protection defends against injection and manipulation, and anti-replay mechanisms with sequence validation mitigate replay and reordering for both payload packets and semantic-related control messages. Robustness to selective dropping can be improved through cross-packet redundancy and unequal error protection for task-critical packets. In addition, receiver-side plausibility checks can detect inconsistencies in recovered semantic information under missing or altered packets, particularly when high-impact packets are selectively removed. Fragmentation and reassembly should be validated and hardened. This prevents malformed fragments and ambiguous reassembly outcomes from corrupting semantic information or bypassing sequence validation and integrity checks. Protocol operations such as acknowledgments, retransmissions, and rate adaptation should be protected against forged feedback. This can be achieved by authenticating feedback messages and applying conservative fallback rules when the reported state is inconsistent or statistically implausible. Configuration and synchronization exchanges should be authenticated, including coding and modulation settings, semantic model versions, codebooks, and knowledge base updates. Strict version validation and rollback prevention are also needed. These measures help avoid persistent transmitter-receiver mismatch that can degrade semantic decoding quality over multiple packets. While these mechanisms improve robustness, padding and redundancy may increase transmission delay. Protecting acknowledgments, retransmissions, and configuration exchanges also adds complexity to protocol control and state management.